OpenStudy (anonymous):

solve for x 1. 4=3sqx^2 2. x^7=1267 3. 81=[2x]^2/3 4. x^4/3

jhonyy9 (jhonyy9):

1. sqrt x^2 =x so than 4=3x --- x=4/3 2. x^7 = 1267 --- x=7^root(1267) 3. 81=cuberoot(4x^2) --- 81^3 = 4x^2 --- x_1,2=+/- sqrt((81^3)/4) = =+/- sqrt(((9^2)^3)/(2^2) =+/- sqrt(((9^3)/2)^2) =+/- (9^3)/2 4. x^(4/3) =cuberoot(x^4) = xcuberoot(x)

OpenStudy (anonymous):

Thank you

jhonyy9 (jhonyy9):

my pleasure good luck bye